OpenStudy (osanseviero):

Mathematical challenge. Today i am going to present the theorem of the identity of a number and its double. It says that any real number is equal to its double... 1 = 2 10 = 20 How? Right now, in an answer, i am going to show the process for do it. Obviously, this is invented by my math teacher. Your challenge will be to find the mistake in the process i am going to put now as an answer Mathematical challenge. Today i am going to present the theorem of the identity of a number and its double. It says that any real number is equal to its double... 1 = 2 10 = 20 How? Right now, in an answer, i am going to show the process for do it. Obviously, this is invented by my math teacher. Your challenge will be to find the mistake in the process i am going to put now as an answer @Mathematics

OpenStudy (anonymous):

these are fun :)

OpenStudy (osanseviero):

I. We put two variables of the same value. \[A = 1\] \[B = 1\] Then, \[A = B\] II. We multiply both sides by A (remember that if you do something to one side, you do the same to the other side). \[A ^{2} = AB\] III. We subtract \[B ^{2}\] to each side. \[A ^{2}-B ^{2} = AB - B ^{2}\] IV. We factorizze \[(a+b)(a-b) = b(a-b)\] V. We divide each side by a-b \[(a+b)(a-b)\div a-b = b(a-b) \div a - b\] VI. Answer A + B = B 2 = 1 Find the mistake!!

OpenStudy (anonymous):

im sure people will find it, this ones very common. :)

OpenStudy (anonymous):

can't divide by a-b because that's 0

OpenStudy (osanseviero):

yep...that was fast xD

OpenStudy (anonymous):

heres a good calculus one.

OpenStudy (anonymous):

what went wrong o.O *gasp*

OpenStudy (anonymous):

\[\sum_{1}^{x} 1 \neq x\] ???

OpenStudy (anonymous):

why doesnt it? I add 3 ones I get 3. I add 5 ones I get 5. I add x 100 ones I get 100. So why wouldnt it be true that when I add x ones I get x? (you are on the right path btw, theres a more concrete way of saying what you mean though, good job!)

OpenStudy (unklerhaukus):

V. is the mistake you cannot divide by a-b because a-b is zero

OpenStudy (anonymous):

oops, "if i add 100 ones I get 100", typo on that "x" >.<

OpenStudy (osanseviero):

Yes UnkleRhaukus...now we are talking of another challenge

OpenStudy (anonymous):

I don't have a clue! Hahaha, help me out!

OpenStudy (anonymous):

in order to take the derivative of a function, the function must be continuous over the domain of integration. The function:\[\underbrace{x+x+\cdots+x}_{x \mbox{\times}}=x^2\] does hold for integers....but what about fractions? or irrational numbers? does it makes sense to add "4/5" 4/5 times? or square roots of 2 square root of 2 times? nope! So this function is only defined on positive integers. Its not continuous, so you cant take the derivative.

OpenStudy (anonymous):

...i dont know why i said integration... i meant over the real numbers.

OpenStudy (anonymous):

im tired >.< my bad.

OpenStudy (anonymous):

Those dumb rules! >.< I'm gonna show everyone :)